A closed gas cylinder contains 0.500 mole H2, 0.300 mole O2, 1.200 mole CO2 at 25 °C and pressures 2.00 atm. Calculate the volume of the cylinder. Then, calculate the partial pressure of O2 inside the cylinder in question.

((R=0.082 L atm mol^-1 K^-1)
1
Expert's answer
2020-06-23T07:16:03-0400

Solution.

"n_{H_2}=0.500mol;"

"n_{O_2}=0.300mol;"

"n_{CO_2}=1.200mol;"

"T=25^oC=298K;"

"P_{total}=2 atm;"

"P_{total}V=n_{total}RT;"

"V=\dfrac{n_{total}RT}{P_{total}};"

"n_{total}=n_{H_2}+n_{O_2}+n_{CO_2};"

"V=\dfrac{(0.500+0.300+1.200)mol\sdot0.082L \sdot atm \sdot mol^{-1}K^{-1}\sdot298K}{2atm}=24.44L;"

"P_{O_2}=n_{O_2}\dfrac{RT}{V};"

"P_{O_2}=0.300mol \dfrac{0.082L\sdot atm\sdot mol^{-1}\sdot K^{-1}\sdot298K}{24.44L}=0.3atm;"

Answer: "V=24.44L;"

"P_{O_2}=0.3atm;"